# LakeVilla: Multi-Table Transactions for Lakehouses

### v1

@@ Abstract.
Data lakehouses (LHs) are at the core of current cloud analytics stacks by providing elastic, relational compute on data in cloud data lakes across vendors. For relational semantics, they rely on open table formats (OTFs). Unfortunately, they have many missing features inherent to their metadata designs, like no support for multi-table transactions and recovery in case of an abort in concurrent, multi-query workloads. This, in turn, can lead to non-repeatable reads, stale data, and high costs in productive cloud systems. In this work, we introduce LakeVilla, a complementary solution that introduces recovery, multi-query/table transactions, and transaction isolation to state-of-the-art OTFs like Apache Iceberg and Delta Lake tables. We investigate its transactional guarantees and show it has minimal impact on performance (2% YCSB writes, 2.5% TPC-DS reads) and provides concurrency control for multiple readers and writers for arbitrary long transactions in OTFs in a non-invasive way.
Analytics, Lakehouse, Object Store, Open Table Formats, Delta Lake, Iceberg, Concurrency, Transactions
copyright: acmcopyrightjournalyear: 2025doi: 10.1145/1122445.1122456conference: We do not know yet; …, 2025; Virtualbooktitle: Booktitleprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Lakehouses for analytical workloads. The past few years have seen the rise of a new type of data management system, called lakehouse (LH), which combines the benefits of low-cost, open-format data lakes and transactional data warehouses (Zaharia et al., 2021; Errami et al., 2023; Levandoski et al., 2024; Kraft et al., 2023; May et al., 2025). With the data gravity leading towards data lakes, LHs for analytical query processing became crucial for various tasks (Zaharia et al., 2021; Inc., 2023; Begoli et al., 2021; Hambardzumyan, 2023; Armenatzoglou et al., 2022; Services, 2020; Kuschewski et al., 2023; Levandoski et al., 2024; Schmidt et al., 2024; Haas and Leis, 2023). LHs use open table formats (OTFs) such as Delta Lake tables (Projects, 2024a), Apache Iceberg (Foundation, 2024a), or Apache Hudi (Foundation, 2021b) to define separate metadata layers on the object store for each table (Zaharia et al., 2021; Armbrust et al., 2020). Those formats implement transactions for single tables and operations, indexes, and time travel (Zaharia et al., 2021; Armbrust et al., 2020), but also lack support for critical features (Armbrust et al., 2020; Jain, 2024).
Missing Features in OTFs. This paper addresses the problem of how to support multi-query and multi-table transactions for LHs while providing strong ACID guarantees. More specifically, we provide native support for the following three missing features: First, the conflict resolution of OTFs cannot track concurrent changes, often resulting in multiple aborts and higher transaction costs. We refer to this feature as (i) Recovery. Second, OTFs today do not support cross-operation dependencies, thereby only allowing simple business logic of single query transactions(Zaharia et al., 2021; Armbrust et al., 2020; Projects, 2024b; Foundation, 2024g; Levandoski et al., 2024). We refer to it as (ii) Complex Transactions. Third, OTFs treat tables separately, leading to inconsistencies and dirty reads under concurrency, and we refer to it as (iii) Isolation. Unfortunately, the lack of support for these features prohibits concurrent writes in a system and can not guarantee properties like referential integrity across tables for concurrent reads. As a result, one cannot execute multi-query and multi-table transactions, forcing use cases for elastic compute or similar (Zaharia et al., 2010; Foundation, 2018; Sethi et al., 2019; Foundation, 2024h) to rely on synchronization methods that block the whole LH (Zaharia et al., 2021; Armbrust et al., 2020). For example, if we normalized our data into different tables like ’students’ and ’lectures’, some concurrent analytics could observe students taking an unknown lecture.
Table 1. Overview of existing formats and extensions and their supported features.
Multi-Query/Table: (i) – (iii)
@@ 1. Introduction
System/Extension Single- query Txn Reco- very (i) Complex Txn (ii) Iso- lation (iii)
[…]
LakeVilla (our approach) $faThumbsOUp \faThumbsOUp \faThumbsOUp$
[…]
@@
State of the Art. Despite multi-query and multi-table transaction support being a highly requested feature in the LH area (Jain and Databricks, 2024; Armbrust et al., 2020; Zaharia et al., 2021), prior work only partially addresses them. In Tab. 1, we summarized the state of the most notable approaches for multi-query and multi-table transaction support in LHs. In summary, none of them adapts to concurrent changes (cf. (i); Recovery) (Foundation, 2024e; Projects, 2024b; Jain et al., 2023; Foundation, 2024f) and some of them currently develop solutions for cross-operation dependency detection (ii) and global transaction isolation (iii). Despite known techniques like the SAGA pattern (Garcia-Molina and Salem, 1987) or MVCC (Bernstein and Goodman, 1983; Neumann et al., 2015; Bernstein et al., 2011; Liu et al., 2020; Iwabuchi et al., 2016) that ensure similar features in traditional database systems, all approaches listed in Tab. 1, add auxiliary components external to the object store but specific to a single OTF format (Jain and Databricks, 2024; licence, 2024; XTable, 2024). In contrast to the original idea of LHs as an open data warehouse (Errami et al., 2023; Zaharia et al., 2021), relying on these external services raises concerns for inter-operability (or unavailability) across vendors, potentially reenabling vendor-lock in. Hence, we aim to support multi-table and multi-query transactions without relying on any auxiliary structures.
@@
LakeVilla (LV) is our extension that implements the three features for (i) recovery, (ii) complex transactions, and (iii) isolation to provide multi-table and multi-query transaction support under concurrency. LV defines a composable set of mechanisms that are non-invasive and compatible with existing OTFs. In contrast to prior work, LV’s mechanisms are integrated into the original protocol and enable each transaction to choose its desired guarantees individually. We ensure LV’s correctness by designing the techniques for the object store that guarantee all transactions result in a valid state of the raw OTF. This ensures a simple integration into existing LH systems and compatibility with clients running the raw OTF. More specifically, LV introduces three modular features: First, LV [Recovery] (LV [R]) uses logical snapshot version markers with transaction sublogs and physical and logical undo and redo information to define a greedy transaction model to support recovery (i). Second, LV [Complex Transactions] (LV [CT]) introduces a marker-based conflict and deadlock detection model among all tables to support complex transactions that include an arbitrary number of queries/operations (ii). Lastly, LV [Isolation] (LV [I]) introduces a global version log that checks global relations among tables and acts as a new layer on the object store above all tables (iii). We also demonstrate how each LV feature strengthens the ACID guarantees of LHs and discuss the performance implications.
@@
Key insights. We show that our LV features guarantee serializability and linearizability for OTFs on the entire LH, not just individual tables or operations. The non-invasive integration of the LV [I] (reads) and LV [R, CT] (writes) can be done with minimal overhead of 2.5% (TPC-DS reads) and 2% (YCSB writes) and incur negligible and predictable costs for typical LH workloads, like analytics.
@@
Our contributions: In this paper, we give an overview of the design area of open table formats in Sect. 2,
@@
    Define non-invasive features LV [R], LV [CT], and LV [I] for (i)–(iii) providing compatibility across all their combinations and the base OTF without vendor lock-in (Sect. 3),
@@
    Show that LV improves the ACID guarantees of OTFs involving multiple tables, queries, and concurrent writes especially guaranteeing Snapshot Isolation up to Serializability (Sect. 4),
@@
    Introduce a LV prototype with minimal overhead and predictable costs when executing our newly adapted YCSB-LH (Cooper et al., 2010), CAB-LH, (van Renen and Leis, 2023), TPC-DS, and microbenchmarks (Sect. 5).
Afterward, we discuss related work in Sect. 6 and conclude in Sect. 7.
2. The Lakehouse Design Area and Structure
@@
Lakehouses (LHs) are an evolution from Data Lakes (Armbrust et al., 2020; Zaharia et al., 2021; Jain et al., 2023) for efficient processing of structured data in object stores using a query engine (Behm et al., 2022; Foundation, 2018, 2024h). LHs aim for comparably lower storage costs in data lakes while maintaining reasonable performance at the expense of data freshness for their expected workloads. The typical use cases of LHs are machine learning and analytical applications with concurrent writes from various data sources with unique characteristics (Zaharia et al., 2021; Armbrust et al., 2020; Foundation, 2024a; Projects, 2024a). This work focuses on analytical applications consisting of an initial load phase and mixed workloads of 40% continuous writes and 60% analytical reads. Writes are typically split between 80% inserts and 20% updates (Schmidt et al., 2024). Across SAP use cases from Ariba (SE, 2024c), S4/HANA (SE, 2024a), and SAP Sales Cloud (SE, 2024b), we found initial load sizes of multiple GBs up to x1 TBs. Workloads have ingestion rates of x103 kB to MB-sized events per hour and complex analytics with up to x102 to x103 concurrent users. The required data freshness lies between a few seconds to daily.
@@
Open table formats (OTFs) and missing features. OTFs on the object store are the basis of LHs (Projects, 2024a; Foundation, 2024a, 2021b). Those formats add a separate metadata layer for each table, organize data, and control client concurrency. This enables advanced analytical features, easy adaption, and builds on open-source formats (Armbrust et al., 2020; Zaharia et al., 2021; Jain et al., 2023; Projects, 2024a; Foundation, 2024a, 2021b). Typically, OTFs use a file hierarchy and rely on atomic requests to the object store (Jain et al., 2023; Begoli et al., 2021; Zaharia et al., 2021; Chen et al., 2022; Armbrust et al., 2020). Two well-known OTFs (Zaharia et al., 2021) (i. e., Snowflake (Snowflake, 2024), Amazon EMR (Goswamia et al., 2022), BigQuery (Google, 2024; Levandoski et al., 2024)) are Delta Lake and Iceberg (Projects, 2024a; Foundation, 2024a). Delta Lake generates log files and checkpoints with unique ascending identifiers (Armbrust et al., 2020; Projects, 2024b). In contrast, Iceberg uses a tree structure to define a current table version represented by a unique branch (Suminto, 2023; Foundation, 2024g). However, those designs and other OTFs share the missing features acknowledged in prior work (e. g., (Zaharia et al., 2021; Armbrust et al., 2020; Jain and Databricks, 2024)) of which we cover the following:
  • (i)
    Recovery: Client crashes and aborts cause unreferenced files, additional costs, high latency, and redundant work.
  • (ii)
    Complex transactions: OTFs do not track dependencies across tables and operations, risking consistency and isolation for arbitrary long transactions.
  • (iii)
    Isolation: Reads might observe dirty data because each table handles commits and version assignments separately.

### v2

# Munich, Germany

@@ Abstract
Data lakehouses (LHs) are at the core of current cloud analytics stacks by providing elastic, relational compute on data in cloud data lakes across vendors. For relational semantics, they rely on open table formats (OTFs). Unfortunately, they have many missing features inherent to their metadata designs, like no support for multi-table transactions and recovery in case of an abort in concurrent, multi-query workloads. This, in turn, can lead to non-repeatable reads, stale data, and high costs in production cloud systems. In this work, we introduce LakeVilla, a modular toolbox that introduces recovery, complex transactions, and transaction isolation to state-of-the-art OTFs like Apache Iceberg and Delta Lake tables. We investigate its transactional guarantees and show it has minimal impact on performance (2% YCSB writes, 2.5% TPC-DS reads) and provides concurrency control for multiple readers and writers for arbitrary long transactions in OTFs in a non-invasive way.
1 Introduction
Table 1: Existing approaches with their supported features and the respective design area
Multi-Query/Table: (i) – (iii) Design Decisions
@@ 1 Introduction
System/Extension Single- query Txn Reco- very (i) Complex Txn (ii) Isola- tion (iii) Built into the OTF structure (D1) Modular solution for OTF transactions (D2)
Apache Iceberg Iceberg_website - , 3 , , ,
Delta Lake journals/pvldb/ArmbrustDPXZ0YM20 ; delta4_multi_table - , 3 3 , ,
Apache Hudi hudi_concurrency_docs - , , , , ,
Project Nessie project_nessie_web - , 3 3 , ,
XTable xtable_web - , , , , ,
DuckLake ducklake - ? - - , ,
LakeVilla (this paper) - - - - - -
Provided: -    Limited/Planned: 3    Unknown due to technical limitations (Section 7): ?    Not provided: ,
Lakehouses for analytical workloads. The past few years have seen the rise of a new type of data management system, called lakehouse (LH), which combines the benefits of low-cost, open-format data lakes and transactional data warehouses journals/jpdc/ErramiHKB23 ; journals/pvldb/Kraft0ZBSYZ23 ; conf/sigmod/LevandoskiCDDEH24 ; conf/cidr/Zaharia0XA21 . With the data gravity leading towards data lakes, LHs for analytical query processing became crucial for various tasks redshift_dl_support ; conf/sigmod/ArmenatzoglouBB22 ; conf/bigdataconf/BegoliGK21 ; journals/pvldb/HaasL23 ; conf/cidr/Hambardzumyan23 ; gartner_cloud ; journals/pacmmod/KuschewskiSAL23 ; conf/sigmod/LevandoskiCDDEH24 ; DBLP:conf/sigmod/SchmidtKHSK24 ; conf/cidr/Zaharia0XA21 . LHs use open table formats (OTFs) such as Delta Lake delta_lake_web , Apache Iceberg Iceberg_website , or Apache Hudi hudi_web to define separate metadata layers on the object store for each table journals/pvldb/ArmbrustDPXZ0YM20 ; conf/cidr/Zaharia0XA21 . Those formats implement transactions for single tables and operations, indexes, and time travel journals/pvldb/ArmbrustDPXZ0YM20 ; conf/cidr/Zaharia0XA21 , but also lack support for critical features journals/pvldb/ArmbrustDPXZ0YM20 ; delta_issues .
Missing Features in OTFs. This paper addresses the problem of how to support multi-query and multi-table transactions for LHs while providing strong ACID guarantees prominently discussed in the community delta4_multi_table ; ducklake . More specifically, we provide native support for the following three missing features: (i) Recovery, (ii) Complex Transactions, and (iii) Isolation. First, the conflict resolution of OTFs cannot track concurrent changes, often resulting in multiple aborts and higher transaction costs delta_issues . We refer to this feature as (i) Recovery. Second, OTFs today do not support cross-operation dependencies, thereby only allowing simple business logic of single query transactions journals/pvldb/ArmbrustDPXZ0YM20 ; conf/sigmod/LevandoskiCDDEH24 ; iceberg_specs ; delta_doc ; conf/cidr/Zaharia0XA21 . We refer to it as (ii) Complex Transactions. Third, OTFs treat tables separately, leading to inconsistencies and dirty reads under concurrency journals/pvldb/ArmbrustDPXZ0YM20 , and we refer to it as (iii) Isolation. Unfortunately, the lack of support for these features prohibits concurrent writes in a system and can not guarantee properties like referential integrity across tables for concurrent reads. As a result, one cannot execute multi-query and multi-table transactions, forcing use cases for elastic compute or similar conf/icde/SethiTSPXSYJHSB19 ; spark_web ; trino_web ; conf/hotcloud/ZahariaCFSS10 to rely on synchronization methods that block the whole LH journals/pvldb/ArmbrustDPXZ0YM20 ; conf/cidr/Zaharia0XA21 . For example, concurrent analytics could observe a reference to a bank account that does not exist yet (see our running example in Section 2).
@@
State of the Art. Multi-query and multi-table transaction support is a highly requested feature for LHs journals/pvldb/ArmbrustDPXZ0YM20 ; delta4_multi_table ; ducklake ; conf/cidr/Zaharia0XA21 that recently got attention from DuckDB ducklake and Databricks delta4_multi_table . In Table 1, we summarize the most notable approaches for multi-query and multi-table transaction support in LHs and their specific design decisions. In summary, most systems do not adapt to concurrent changes (cf. (i); Recovery) conf/cidr/0001KPDSZ23 ; spark_doc ; ib_reliability_docs ; delta_doc . At the same time, solutions for cross-operation dependency detection (ii) and global transaction isolation (iii) are either not present or currently in development for some systems. Despite known techniques like the SAGA pattern conf/sigmod/Garcia-MolinaS87 or MVCC journals/tods/BernsteinG83 ; journals/pvldb/BernsteinRWY11 ; conf/ipps/IwabuchiSPEGM16 ; journals/pvldb/LiuHMCLSSSAA20 ; conf/sigmod/0001MK15 that ensure similar features in traditional database systems, all previous approaches listed in Table 1, add auxiliary components external to the object store but specific to a single OTF format project_nessie_web ; xtable_web ; delta4_multi_table ; ducklake (D1). At the time of writing, only DuckLake ducklake and Delta Lake 4.0 delta4_multi_table support multi-query and multi-table transactions in OTFs, but they diverge significantly from OTF principles by offloading all metadata to DuckDB ducklake and/or heavily modifying the OTF structure delta4_multi_table . Additionally, they are facing technical limitations, making their position in Table 1 unclear (see Section 7). However, relying on these external services raises concerns for interoperability across vendors, potentially reenabling vendor-lock in, contradicting the original idea of LHs as an open data warehouse journals/jpdc/ErramiHKB23 ; conf/cidr/Zaharia0XA21 . Hence, we aim to support multi-table and multi-query transactions without any external structures required.
@@
LakeVilla (LV) is our toolbox that implements the three features (i) recovery, (ii) complex transactions, and (iii) isolation to provide multi-table and multi-query transaction support under concurrency. LV defines a modular set of modules that are non-invasive and compatible with existing OTFs (D2). In contrast to prior work project_nessie_web ; xtable_web ; journals/pvldb/ArmbrustDPXZ0YM20 ; ducklake , LV’s modules #1–#4.2 are integrated into the original OTF protocol and enable each transaction to choose its desired guarantees individually (D1). We ensure LV’s correctness by guaranteeing that modules result in a valid state of the raw OTF. This allows a simple integration into existing LH systems and compatibility with clients running the raw OTF. More specifically, LV composes #1–#4.2 into three modular features: First, LV[Recovery] (LV[R]) uses logical snapshot version markers (#1) with physical and logical transaction sublogs (#2) defining a greedy transaction model to support recovery (i). Second, LV[Complex Transactions] (LV[CT]) introduces a marker-based conflict and deadlock detection model (#3) among all tables to support complex transactions with an arbitrary number of queries/operations (ii). Lastly, LV[Isolation] (LV[I]) introduces a global version log (#4.1 and #4.2) that checks global relations among tables and acts as a new layer on the object store above all tables (iii). We also demonstrate how each LV feature strengthens the ACID guarantees of LHs and discuss its implications.
@@
Key insights. We show that LV guarantees serializability and linearizability for OTFs on the entire LH, not just individual tables or operations. The non-invasive integration of LV[I] (reads) and LV[R, CT] (writes) can be done with minimal overhead of 2.5% (TPC-DS reads) and 2% (YCSB writes) and incur negligible and predictable costs for typical LH workloads, like analytics.
@@
Our contributions: In this paper, we give an overview of OTFs and their design area in Sections 2 and 3,
@@
    Define modules #1–#4.2 and compose them into the features LV[R], LV[CT], and LV[I] for (i)–(iii) providing compatibility across all their combinations and the base OTF without introducing vendor lock-in (Section 2),
@@
    Show that LV improves the ACID guarantees of OTFs involving multiple tables, queries, and concurrent writes, especially guaranteeing Snapshot Isolation up to Serializability (Section 5),
@@
    Introduce an LV prototype with minimal overhead and predictable costs when executing our adapted YCSB-LH conf/cloud/CooperSTRS10 , CAB-LH journals/pvldb/RenenL23 , TPC-C tpcc , LSTBench 2024lstbench , and microbenchmarks (Section 6).
Afterward, we discuss related work and conclude in Sections 7 and 8.
2 Lakehouses & Open Table Formats
@@
Lakehouses (LHs) are an evolution from Data Lakes journals/pvldb/ArmbrustDPXZ0YM20 ; conf/cidr/0001KPDSZ23 ; conf/cidr/Zaharia0XA21 for efficient processing of structured and unstructured data in object stores using a query engine conf/sigmod/BehmPAACDGHJKLL22 ; spark_web ; trino_web . LHs aim for comparably lower storage costs in data lakes while maintaining reasonable performance at the expense of data freshness for their expected workloads. The typical use cases of LHs are machine learning and analytics with concurrent writes from various data sources with unique characteristics journals/pvldb/ArmbrustDPXZ0YM20 ; Iceberg_website ; delta_lake_web ; conf/cidr/Zaharia0XA21 . This work focuses on analytical applications with concurrent writes, typically consisting of an initial load phase and mixed workloads of 15% continuous writes, 30% analytical reads, and 55% read-writes as observed by industry and publicly available datasets (e. g., Snowset snowflake-nsdi20 , Redset Renen2024 ). Writes are typically split between 80% inserts and 20% updates DBLP:conf/sigmod/SchmidtKHSK24 . Across SAP use cases from Ariba sap_ariba , S4/HANA sap_hana , and SAP Sales Cloud sap_sales_cloud , we found initial load sizes of multiple GBs up to single TBs. Workloads have ingestion rates of x103 kB to MB-sized events per hour and complex analytics with up to x102 to x103 concurrent users. The required data freshness lies between a few seconds to daily.
@@
Open table formats (OTFs) on the object store are the basis of LHs hudi_web ; Iceberg_website ; delta_lake_web . Those formats add a separate metadata layer for each table, organize data, and control client concurrency. This enables advanced analytical features, easy adaptation, and builds on open-source formats journals/pvldb/ArmbrustDPXZ0YM20 ; conf/cidr/0001KPDSZ23 ; hudi_web ; Iceberg_website ; delta_lake_web ; conf/cidr/Zaharia0XA21 . Typically, OTFs use a file hierarchy and rely on atomic requests to the object store journals/pvldb/ArmbrustDPXZ0YM20 ; conf/bigdataconf/BegoliGK21 ; conf/cbd/ChenSLLJ22 ; conf/cidr/0001KPDSZ23 ; conf/cidr/Zaharia0XA21 . Two well-known OTFs conf/cidr/Zaharia0XA21 (i. e., Snowflake snowflake_iceberg_web , Amazon EMR lakehouse_aws_emr , BigQuery delta_lake_bigQuery ; conf/sigmod/LevandoskiCDDEH24 ) are Delta Lake and Iceberg Iceberg_website ; delta_lake_web .
The metadata layers of OTFs are essential and the biggest difference from previous architectures conf/cidr/Zaharia0XA21 . LV operates directly within the OTFs used by LHs. For simplicity, this paper uses Delta Lake to explain LV. However, LV can be analogously applied to Apache Iceberg Iceberg_website or other OTFs with minimal changes.
Figure 1: Metadata Layer design of Delta Lake in a banking example (running example) (draw.io)
Delta Lake delta_lake_web is one of the most well-known and practically used OTFs conf/cidr/Zaharia0XA21 (i. e., Amazon EMR lakehouse_aws_emr , BigQuery delta_lake_bigQuery ; conf/sigmod/LevandoskiCDDEH24 ). The structure of a delta lake table delta_lake_web ; delta_doc is depicted in Figure 1 for a fictional table Bank x. A Delta Lake table consists of a metadata layer structured like a linked list and a data layer. A snapshot is defined by its respective log file and encodes all file operations to advance the previous snapshot to its own journals/pvldb/ArmbrustDPXZ0YM20 . These operations range from adding or removing files to adding metadata, and more journals/pvldb/ArmbrustDPXZ0YM20 . To retrieve the full table version of a snapshot, engines must access the respective logs. Figure 1 defines snapshots 0, 1, and 2 with logs (0), (1), and (2). Table version 0 is fully defined by Log (0) and contains Files #1 and #2, while version 1 is defined by Logs (0) and (1) and consists of Files #2 and #3. To avoid excessive sequential accesses to all logs, Delta Lake implements a pointer to the last checkpoint created, usually stored in a separate file, and regular checkpoints stored in Parquet format journals/pvldb/ArmbrustDPXZ0YM20 . A checkpoint summarizes all valid table instructions and snapshots at the time of its creation journals/pvldb/ArmbrustDPXZ0YM20 . In Figure 1, the last checkpoint contains the instructions of logs (0) and (1). Hence, a client only reads the checkpoint and Log (2) to retrieve version 2.
2.1 Running Example: Delta Lake Operations
Reads: Check. We showcase the operations within a Delta Lake table using a simple banking example shown in Figure 1. All accounts of bank x are stored within the table and contain their id and balance. First, the bank retrieves a check request for account 2. Hence, the used engine retrieves the current snapshot by issuing a LIST request on the metadata directory and determines the required logs journals/pvldb/ArmbrustDPXZ0YM20 . For snapshot (0), it only reads Log (0) and accesses Files #1 and #2. After the required data is retrieved, the engine can commit without changing the metadata layer.
Writes: Deposit and Withdraw. The first write request the bank retrieves is a deposit request for account 1 of 50\$. The engine starts exactly like a read request by determining the newest snapshot, reading the respective logs, and assembling the current version. However, since files can only be overwritten in data lakes, it continues to create a new file #3 with the updated balance of 5050\$ for account #1. When the operation is committed, the engine adds log (1), removing file #1 with the outdated value and adding #3. This effectively creates snapshot (1). Due to the engine settings, it creates a checkpoint containing the information of all current logs (0 and 1) and sets the last_checkpoint pointer.
Next, 200\$ is withdrawn from Acc. 2. Similar to before, an engine receives snapshot (1) but reads the last checkpoint instead of Log (0) and (1). The required file #2 is used as the basis for the updated balance 4800\$ written as file #4. During commit, the engine creates snapshot (2) by adding Log (2), replacing File #2 with #4. Here, the engine decides not to create a checkpoint and finishes its operation.
3 The Lakehouse Design Area
3.1 Missing Features of OTFs
Despite different metadata layers, all OTFs share similar missing features by default, as also acknowledged in prior work journals/pvldb/ArmbrustDPXZ0YM20 ; delta4_multi_table ; conf/cidr/Zaharia0XA21 . In this paper, we cover:
(i) Recovery: No Duplicated Work on Aborts. All OTFs define a single table-level synchronization point in their metadata layers in the form of appending logs or branches to enable optimistic concurrency control journals/pvldb/ArmbrustDPXZ0YM20 ; iceberg_specs . Hence, a transaction can only be committed if it successfully passes this bottleneck. All transactions based on the same accessed branch/log must abort as the new commit might have accessed the same data iceberg_specs . However, OTFs do not provide any means to recover or undo a failed transaction. This can result in unreferenced data files, additional costs, and high latency. Redoing the transaction results in redundant work in the form of additional requests to the object store, especially uploading duplicate data files. The impact of this missing feature can intensify for more complex operations, as they might starve or fail with the standard OTF concurrency control, as seen in section 6 for TPC-C.
(ii) Complex Transactions: Concurrent Multi–query and Multi–table Transactions. The OTF design decision to focus on transactions consisting of a single SQL query (single-query) and accessing strictly one table (single-table) led to publishing each (sub-) operation after its execution (a) and a table-level design (b) of the metadata journals/pvldb/ArmbrustDPXZ0YM20 ; iceberg_specs ; conf/cidr/0001KPDSZ23 ; delta_doc . (a) breaks the ACID guarantees for multi-query transactions: When simply applying the current designs, intermediate changes of multi-query transactions would become immediately visible (violating atomicity) and could get overwritten by other concurrent transactions (violating consistency and isolation). Further, (b) prevents multi-table transactions: An abort in one table could lead to a partially committed transaction (violating consistency) that can not be undone without potentially aborting newer transactions or deleting committed values (violating durability and introducing a risk of cascading aborts). Thus, users must either break isolation or risk starvation of complex queries for complex transactions.
(iii) Isolation: Controlling global versions. Due to the independent metadata layers in OTFs, multi-table transactions must interact with the metadata layer of each table separately Iceberg_website ; iceberg_specs ; delta_lake_web ; delta_doc . This design decision ignores any global information about the tables’ states and dependencies, assigning versions on the table level. Under concurrency, other transactions might observe an inconsistent combination of table versions (dirty read), as some tables might adopt changes of another transaction earlier than others. Most prominently, LHs cannot prevent read transactions from observing potentially partially committed multi-table transactions even with the restriction to a single writer conf/cidr/Zaharia0XA21 . A global lock, as suggested by journals/pvldb/ArmbrustDPXZ0YM20 ; iceberg_specs ; delta_doc , can solve the problem but reduces the LH’s throughput due to partially enforcing sequential execution. As concurrent reads and read-writes are the dominant workload of LHs, it becomes essential to guarantee transaction isolation of read and write transactions.
3.2 Design Questions